Circular No. 6718

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970815 D. A. Smith, A. M. Levine, E. H. Morgan, and A. Wood report, on be-

half of the RXTE/ASM team at the Massachusetts Institute of Technology and Goddard Space Flight Center, their independent detection with the ASM detector of an x-ray burst that they identify as a counterpart to the γ-ray burst source detected by the Compton Gamma Ray Observatory’s BATSE detector on Aug. 15.50491 UT (posted at http://lheawww.gsfc.

nasa.gov/docs/gamcosray/legr/bacodine/baco notes.html): “The

burst rise was detected in one camera on Aug. 15.50597. At Aug. 15.50623, the cameras were rotated through 6^{o}, such that the burst came into the field-of-view of a second camera. Observations with two cameras at different orientations enabled us to estimate the position of the burst source to be within an error box having the following corners: α = 16^{h}06^{m}36^{s}, δ = +81^{o}27′36′′; 16^{h}10^{m}50^{s}, +81^{o}33′36′′; 16^{h}08^{m}10^{s}, +81^{o}31′12′′; 16^{h}09^{m}17^{s}, +81^{o}30′00′′ (equinox 2000.0). The burst lasted ∼ 130 s and showed a double-peaked structure, reaching a maximum intensity of almost 2 Crab (2–12 keV) in the second peak.”

RR TELESCOPII P. Whitelock, South African Astronomical Observatory (SAAO), re-

ports that the symbiotic Mira-type variable RR Tel is > 1 mag fainter in the J band than is usual at this phase of its pulsation cycle — as indicated by the following J magnitudes from F. Marang at the SAAO 0.75-m telescope in Sutherland: May 22, 8.0; Aug. 12, 7.3. The star’s colors are red, suggesting increased dust obscuration.

V4334 SAGITTARII E. F. Guinan, G. McCook, and S. Margheim, Villanova University,

write: “UBVRI photoelectric photometry of the final helium-flash candidate V4334 Sgr (cf. IAUC 6322, 6323, 6325, 6328) has been conducted with the Four College Consortium 0.8-m APT in Arizona over 23 nights during June and July. The photometry reveals low-amplitude (0.07 mag in V) semi-periodic light variations having a characteristic timescale of ∼ 22 ± 3 days. Smaller-amplitude (0.02 mag in V) light variations also are present with timescales of 6–10 days. Light variations like these are commonplace for many post-AGB stars and most likely arise from pulsations. The mean magnitude during this interval was V ∼ 11.1, similar to the brightness that the star has maintained since its rapid brightening.”

1997 August 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green